a rectangle is 2 feet long and 1.5 feet wide,, what is the distance if you measure it diagonally?
Found 3 solutions by Alan3354, solver91311, checkley77:
Answer by Alan3354(69443)
Answer by solver91311(24713)



Use Pythagoras:



so



Or in your case:



You can do your own arithmetic.

John


Answer by checkley77(12844)

These 2 sides 2 ft. & 1.5 ft. along with the diagonal of the rectangle form a right triangle.
Thus we have the formula for the 3 sides of a right triangle:
a^2+b^2=c^2
2^2+1.5^2=c^2
4+2.25=c^2
c^2=6.25
c=sqrt6.25
c=2.5 ft. ans. for the length of the long side (hypotenuse) of the triangle and the diagonal of the rectangle.
Proof:
2^2+1.5^2=2.5^2
4+2.25=6.25
6.25=6.25